A store is selling CD's for $10.50 and for $8.50. You buy 10 total and spend $93.00. How many of each price did you buy?
Answer by Earlsdon(6294)

Let x = the number of CDs at $10.50 each, and y = the number of CDs at $8.50 each.
The total number of CDs is: x + y = 10 Rewrite this as: x = 10 - y and substitute in the equation below:
The total cost of CDs is: (x)($10.50) + y($8.50) = $93.00
(10 - y)(10.5) + y(8.5) = 93. Simplify and solve for y.
105 - (10.5)y + (8.5)y = 93 Subtract 105 from both sides.
-10.5y + 8.5y = - 12
-2y = -12 Divide both sides by -2.
y = 6
x = 10 - y
x = 10 - 6
x = 4
Check:
4($10.50) + 6($8.50) = $42.00 + $51.00 = $93.00